OpenStudy (anonymous):

multiply and simplify sqrt{?}5b^5 sqrt{?}15c^6??????

OpenStudy (anonymous):

\[\sqrt{5}*b^{5}*\sqrt{15}*c^6=\sqrt{(25*3)}*b^5c^6=5b^5c^6\sqrt{3}\]\[\sqrt{5b^5}*\sqrt{15c^6} = b^2c^3\sqrt{75b}=5b^2c^3\sqrt{3b}\] I wasn't completely sure what you wanted answering, so I answered the two things I took that to mean. I hope they are of help.